# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v2

@@ 2.5 AI Model Taxonomy
  • Advanced Deep Learning (A-DL) Models: Incorporates more complex or novel architectures, advanced training techniques, and elements of other AI fields like reinforcement learning into classic deep networks. Examples include graph convolutional networks (GCN) [96], Transformers [97], Generative Adversarial Networks (GANs) [98], etc.
@@ 3.1 Study Selection
We adopt the Preferred Reporting Items for Systematic Reviews and Meta-Analysis (PRISMA) [99] to conduct our survey, and the flow diagram of PRISMA is provided in the appendices. We summarize the literature selection process as follows: During the paper identification stage, we retrieve 840 records from PubMed, 895 records from WoS, and 596 records from Google Scholar. By comparing the collected records across these three databases, a total of 2069 records remain after 262 duplicate records are removed. In the paper screening stage, after screening the titles and abstracts of the remaining 2069 papers according to the inclusion criteria, 243 articles are further selected for the full-text assessment. Finally, after an in-depth review of the full texts of these selected papers, 103 articles are excluded and 140 regular articles are included in this survey for further information extraction and content summarization. The list of excluded papers from the full-text screening stage, along with the reasons for their exclusion, is detailed in our appendices. Moreover, 14 eligible survey papers [46, 34, 35, 47, 36, 37, 38, 39, 100, 40, 41, 42, 43, 44] are found in the screening stage. From these survey papers, we identify another 99 potentially-matched articles via screening their titles and abstracts. After the full-text assessment, we include 24 regular articles from those. In summary, we include a total of 164 eligible articles, comprising 140 from the academic databases and 24 from survey papers, in our survey.
@@ 3.2 Statistics of Selected Papers
To provide an intuitive overview for all 164 selected papers, we present the distribution of papers from different aspects: year of publication (see Fig. 2), country of publication, type of NDs studied, type of gait data, type of AI models, and size of samples (see Fig. 8).
@@ 3.2.2 Distribution of Countries
The results in Fig. 8f suggest that China, India, and the United States (US) are the top 3 contributing countries to this research field, with 29 papers ($17.7\%$), 15 papers ($9.1\%$) each from India and the US, respectively. This approximately matches the real-world situation where Chinese and Indian NDs patients constitute the two largest groups globally, thus a broader research resources and efforts are put into this important topic. It is also observed that, within the top 10 countries with the most publications, the number of developed countries (US, UK, Italy, Canada, Australia) participated in this research field is equal to that of the developing countries (China, India, Colombia, Iran, Brazil). However, among these 10 countries, the total publications from developing countries constitutes $59.6\%$ (68 publications), whereas that from developed countries accounts for only $40.4\%$ (46 publications).
@@ 3.2.3 Distribution of ND Types
Fig. 8b shows the numbers of papers focused on different NDs types. Note that the studies containing diagnosis of multiple NDs (19 articles) are also included into each NDs type ($e.g.$, a study simultaneously diagnosing PD, HD, and ALS are counted as one PD/HD/ALS study). The statistics show that PD receives the most research focus among all the NDs, including a total of 143 articles ($87.2\%$). The second most studied ND’s type is HD with 26 papers ($15.9\%$), while there are 22 ($13.4\%$), 6 ($3.7\%$), 5 ($3.0\%$) papers focused on ALS, MS, and AD respectively. Interestingly, although AD is the most populated diseases among all NDs (see Fig. 2), the study of its diagnosis based on gait has yet to receive much attention, possibly because other NDs patients, such as PD patients, may exhibit more evident abnormal gaits that can be exploited to perform more reliable diagnosis. Promisingly, more and more studies [48, 101, 102, 77, 103, 104, 105, 106] have demonstrated the feasibility of leveraging gait data to identify HD, AD, ALS, and MS, therefore a deeper exploration of these research directions should be conducted.
@@ 3.2.4 Distribution of Gait Data Types
As reported in Fig. 8d, 137 studies ($83.5\%$) utilize sensors including Inertial Measurement Units (IMUs) and force/pressure sensitive resistors (FSRs) to collect vertical Ground Reaction force (vGRF) data. These sensors can capture detailed and dynamic information about body movements and forces involved in walking to analyze gait cycles and parameters. There are 13 studies ($7.9\%$) focused on vision-based devices such as RGB cameras and infrared cameras to extract gait patterns from videos. A same number of studies (13 articles) combine both sensor-based and vision-based devices to collect multi-modal gait data for learning.
Fig. 8e further shows the distribution of sensor types used in 137 sensor-based studies: The force sensors (FS) ($e.g.,$ force/pressure sensitive resistors) and IMUs are the two most widely-used sensor types, occupying $56.2\%$ (77 studies) and $31.4\%$ (43 studies) respectively. For other sensors (17 studies), there are 5 studies combining both FS and IMUs as gait data sources, while 5 studies utilize gait analysis system (GAS) ($e.g.,$ mats, walkways, treadmills). The low-cost and non-invasive motion capture device, Kinect [107, 82], is also used as both a sensor and a vision-based device for gait data collection (2 studies). We provide a further discussion on exploiting sensor-based ($e.g.,$ Kinect-based) 3D skeleton data for NDs diagnosis in Sec. 5.5.
@@ 3.2.5 Distribution of AI Models
As shown in Fig. 8a, in the past decade, only 17 articles ($10.4\%$) propose novel A-DL models for NDs diagnosis from gait, while most of the studies use conventional AI methods such as C-ML and C-DL with 104 ($63.4\%$) and 43 articles ($26.2\%$) respectively. However, most conventional AI models possess limited performance on large-scale medical data, thereby underscoring the challenge and urgency of developing more advanced AI models in this field. Such development requires more efforts and cooperation from AI scientists and healthcare professionals in the future.
@@ 3.2.6 Distribution of Sample Sizes
As shown in Fig. 8c, 137 papers ($83.5\%$) conduct their experiments with more than 30 samples ($i.e.,$ number of patients plus number of controls), and the number of papers that possess small-size samples ($10\leq n\leq 30$) only occupies $16.5\%$ (27 papers). This suggests that most AI models require a relatively larger size of samples and gait data to obtain satisfactory performance, otherwise it might be insufficient for obtaining reliable diagnosis. This also aligns with the recent advancements of medical deep learning based AI models that necessitate massive medical labeled data.
@@ 4.1 Quality Evaluation Criterion
To comprehensively measure the quality of each study, we empirically grade the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples at three levels with the score 1, 2, or 3. We sum their scores and divide them by the maximum score to get a normalized quality score $S_{Q}$ ranging from 0 to 1. $S_{Q}$ is computed for each study, which are defined as follows:
  • Score ($S_{n}$) for Novelty of AI Models
@@
      Score $S_{n}=1$: The main proposed method belongs to or contains only C-ML models.
@@
      Score $S_{n}=2$: The main proposed method belongs to or contains C-DL models.
@@
      Score $S_{n}=3$: The main proposed method belongs to or contains A-DL models.
@@
Table 1: Number of studies related to different NDs (PD, HD, MS, AD, ALS) and different model types (C-ML, C-DL, A-DL). “Combinatorial” denotes studies that simultaneously contain diagnosis of multiple NDs including PD, ALS, and HD. $S_{Q}>0.6$ indicates the number of high-quality studies.
        Type         C-ML         C-DL         A-DL         Total         $\boldsymbol{S_{Q}>0.6}$
        PD         77         34         13         124         56
@@
        11         6         2         19         8
        HD         6         1         0         7         1
        MS         4         1         1         6         2
        AD         3         1         1         5         2
        ALS         3         0         0         3         0
        Total         104         43         17         164         —
        $\boldsymbol{S_{Q}>0.6}$         23         29         17         —         69
@@ 4.2 Parkinson’s Disease (PD)
As shown in Table 2, 3, 4, 5, and 6, there are 124 studies that focus on PD diagnosis based on human gait. It is worth mentioning that there are 19 studies, each of which simultaneously includes the diagnosis of PD and other NDs, as shown in Table 6 (referred to as “Combinatorial Diagnosis of Different NDs”). We separately discuss these studies in Sec. 4.3. Based on our survey, there are 77, 34, and 13 PD studies using C-ML, C-DL, and A-DL models respectively. It should be noted that there are a large number of PD studies (up to 124 studies). Since the studies with high-quality AI models and diagnostic results are the focus of our survey, we first categorize all studies by three AI model types (C-ML, C-DL, A-DL), and then select representative studies with quality score $S_{Q}>0.6$ for content summary and analysis.
@@ 4.2.1 C-ML Models for PD Diagnosis
[108] (Table 2, ID = P-7) utilizes Hidden Markov Models (HMM) with Gaussian Mixtures to classify gait data and differentiate Parkinson’s disease (PD) patients from healthy subjects. Using stride interval as the distinguishing feature, the HMM method achieves an accuracy rate of 90.3%. In [80] (ID = P-10), the authors presents a methodology using Microsoft Kinect sensors to detect gait attributes and recognize gait of PD patients. The leg length, stride length, and gait velocity are extracted to achieve high-accuracy classification using neural networks (NN). This study suggests the potential of using Kinect sensors for gait analysis and disorder recognition. [109] (ID = P-14) proposes a radial basis function neural networks (RBF-NN) based method for classifying PD patients and healthy individuals using gait characteristics derived from vertical ground reaction forces (vGRF), achieving over 90% accuracy. [110] (ID = P-16) focuses on a technique called Shifted One-Dimensional Local Binary Patterns (Shifted 1D-LBP) to track local changes in time series data. Combining with machine learning model, this work demonstrate Shifted 1D-LBP’s potential in real-time NDD detection applications. The study of [57] (ID = P-20) utilizes the vGRF data to extract gait features and feeds them into the SVM classifier to obtain a good accuracy result of 93.6%. [111] (ID = P-21) explores the use of wavelet analysis combined with SVM for the identification of PD based on spatio-temporal gait variables, achieving a classification accuracy of 90.32%. The study demonstrates the potential of using wavelet analysis as an efficient method for classifying PD and healthy subjects. [112] (ID = P-30) utilizes a wearable inertial device to collect motor data from healthy subjects, individuals with idiopathic hyposmia (IH), and PD patients during lower limb assessments, and develops a machine learning-based system with SVM, RF, NB to classify the three groups. In [113] (ID = P-32), the authors quantify gait parameters objectively using wearable accelerometers, and compare them with a motion capture system to automatically discriminate patients with PD using machine learning models such as SVM, KNN, and DT. The proposed approach achieves an accuracy of approximately 89.1% and suggests that the wearable accelerometer-based system is suitable for assessing and monitoring PD and Freezing of gait (FOG) in real-life scenarios. [114] (ID = P-38) emphasizes of using a hybrid of linear and non-linear features extracted from human gait time series data to improve the NDD classification system which employs SVM with a radial basis kernel function. [115] (ID = P-42) leverages different classic machine learning methods (KNN, DT, RF, NB, SVM, K-Means) to identify PD from vGRFs collected from gait cycles. The structure of PD recognition process from data acquisition to performance evaluation is also systematically presented in this study. [116] (ID = P-47) devises an automated system with force sensors to extract gait parameters and then use these parameters as input to the SVM model to diagnosis PD. [117] (ID = P-50) proposes a novel method for classifying gait patterns between patients with PD and healthy controls. This method reconstructs the phase space of vGRF and decomposes the gait dynamics using empirical mode decomposition, which are fed to neural networks (NN) for classification with accuracy of 98.8%. In [118] (ID = P-57), the authors explore the effectiveness of Extreme Gradient Boosting (XGBoost) and Artificial Neural Network (ANN) models for monitoring the gait of Parkinson’s Disease patients, achieving 41.0%(double check the figure) accuracy. The study of [119] (ID = P-60) explores the use of single-sensor accelerometer data from healthy subjects to extract features relevant to multi-sensor accelerometer gait data for Parkinson’s disease classification. They use a pre-trained convolutional autoencoder as the source model and a simple multi-layer perceptron as the target model to classify PD and achieve 68.92% accuracy. A machine learning based gait classification system with different models such as SVM and RT is presented in [120] (ID = P-65) to assist in diagnosing PD form vGRF. [121] (ID = P-71) extracts spatio-temporal gait features such as stride time and stance time from vGRF and exploits different classic machine learning models (KNN, NB, SVM) to classify health control and PD subjects with an accuracy of 88.5-98.8%. In [122] (ID = P-81), a time-varying singular value decomposition method is designed to extract the most useful and informative part of the vGRF signal data, and then sparse non-negative least squares (NNLS) approach is leveraged to classify PD and HCs, achieving 97.2% accuracy. [123] (ID = P-87) utilizes RF to identify PD patients from vGRF and achieves up to 98.8% accuracy. The SVM model is used in [124] (ID = P-111) to classify PD and HC gait from vGRF data, with an accuracy of 98.2%.

### v3

@@ 2.5 AI Model Taxonomy
  • Advanced Deep Learning (A-DL) Models: Incorporates more complex or novel architectures, advanced training techniques, and elements of other AI fields like reinforcement learning into classic deep networks. Examples include graph convolutional networks (GCN) [101], Transformers [102, 103], Generative Adversarial Networks (GANs) [104], etc.
@@ 3.1 Study Selection
We adopt the Preferred Reporting Items for Systematic Reviews and Meta-Analysis (PRISMA) [105] to conduct our survey, and the flow diagram of PRISMA is provided in the appendices. We summarize the literature selection process as follows: During the paper identification stage, we retrieve 845 records from PubMed, 895 records from WoS, and 596 records from Google Scholar. By comparing the collected records across these three databases, a total of 2074 records remain after 262 duplicate records are removed. In the paper screening stage, after screening the titles and abstracts of the remaining 2074 papers according to the inclusion criteria, 248 articles are further selected for the full-text assessment. Finally, after an in-depth review of the full texts of these selected papers, 103 articles are excluded and 145 regular articles are included in this survey for further information extraction and content summarization. The list of excluded papers from the full-text screening stage, along with the reasons for their exclusion, is detailed in our appendices. Moreover, 14 eligible survey papers [46, 34, 35, 47, 36, 37, 38, 39, 106, 40, 41, 42, 43, 44] are found in the screening stage. From these survey papers, we identify another 99 potentially-matched articles via screening their titles and abstracts. After the full-text assessment, we include 24 regular articles from those. In summary, we include a total of 164 eligible articles, comprising 145 from the academic databases and 24 from survey papers, in our survey.
@@ 3.2 Statistics of Selected Papers
To provide an intuitive overview for all 169 selected papers, we present the distribution of papers from different aspects: year of publication (see Fig. 2), country of publication, type of NDs studied, type of gait data, type of AI models, and size of samples (see Fig. 8).
@@ 3.2.1 Distribution of Publication Years
3.2.2 Timeline of Technical Advancements
Figure 9 illustrates the timeline of technical advancements in gait-based PD diagnosis. The timeline, spanning from 2012 to 2023, highlights significant advancements in both the models and data types utilized for gait-based PD diagnosis. Throughout this period, SVM, as exemplified by models in [88] (ID = P-1) and [107] (ID = P-122) , consistently demonstrated their efficacy and reliability in C-ML approaches for diagnosing PD from 2012 to 2023.
In 2016, a landmark innovation emerged with the introduction of neural networks in [108] (ID = P-15), marking the first use of DL for PD diagnosis. By 2023, the field saw the development of a more sophisticated DL model, static-dynamic temporal networks, which integrated one-dimensional and two-dimensional convolutional networks along with attention mechanisms. This advanced model, featured in [109] (ID = P-123), achieved a remarkable diagnostic accuracy of 96.7%.
Sensor-based data, including EMG and IMUs, gained prominence as effective tools for enhancing diagnostic precision from 2012 to 2023, as evidenced by their application in [77] (ID = P-2) and [110] (ID = P-112). Vision-based approaches also evolved, with models like [86] (ID = P-43) and [111] (ID = P-121) utilizing Kinect and smartphone cameras, respectively, to extract spatial and temporal gait features from video data.
In 2022, multimodal methodologies gained traction. For instance, [46] (ID = P-113) combined sensor data from DSLR cameras and passive markers to integrate complementary inputs, significantly improving the robustness of PD diagnosis. The trend toward combining sensor and vision data culminated in [82] (ID = P-106), where advanced DL models leveraged inputs from Kinect cameras, vGRF, and IMUs for a comprehensive analysis of gait dynamics.
This timeline reflects the continuous evolution of gait-based PD diagnostic methods, progressing from traditional machine learning techniques to cutting-edge deep learning models and multimodal approaches, paving the way for more precise and holistic diagnostic capabilities. We also provide a more comprehensive table on the timeline of gait-based PD diagnosis and the combinatorial diagnosis of different NDs in our appendices.
Figure 9: Timeline for technical advancement of gait-based PD diagnosis.
@@ 3.2.3 Distribution of Countries
The results in Fig. 10b suggest that China, the United States (US), and India are the top 3 contributing countries to this research field, with 31 papers ($18.3\%$), 15 papers ($8.9\%$), and 15 papers ($8.9\%$), respectively. This approximately matches the real-world situation where NDs patients of these countries constitute the three largest groups globally, thus a broader research resources and efforts are put into this important topic. It is also observed that, within the top 10 countries with the most publications, the number of developed countries (US, UK, Italy, Canada, Australia) participated in this research field is equal to that of the developing countries (China, India, Colombia, Iran, Brazil). However, among these 10 countries, the total publications from developing countries constitutes $59.8\%$ (70 publications), whereas that from developed countries accounts for only $40.2\%$ (47 publications).
@@ 3.2.4 Distribution of ND Types
Fig. 8b shows the numbers of papers focused on different NDs types. Note that the studies containing diagnosis of multiple NDs ( 21 articles) are also included into each NDs type ($e.g.$, a study simultaneously diagnosing PD, HD, and ALS are counted as one PD/HD/ALS study). The statistics show that PD receives the most research focus among all the NDs, including a total of 147 articles ($69.7\%$). The second most studied ND’s type is HD with 28 papers ($13.3\%$), while there are 24 ($11.4\%$), 6 ($2.8\%$), 6 ($2.8\%$) papers focused on ALS, MS, and AD respectively. Interestingly, although AD is the most populated diseases among all NDs (see Fig. 2), the study of its diagnosis based on gait has yet to receive much attention, possibly because other NDs patients, such as PD patients, may exhibit more evident abnormal gaits that can be exploited to perform more reliable diagnosis. Promisingly, more and more studies [53, 112, 113, 82, 114, 115, 116, 117] have demonstrated the feasibility of leveraging gait data to identify HD, AD, ALS, and MS, therefore a deeper exploration of these research directions should be conducted.
@@ 3.2.5 Distribution of Gait Data Types
As reported in Fig. 8d, 142 studies ($84.0\%$) utilize sensors including Inertial Measurement Units and force/pressure sensitive resistors to collect vertical Ground Reaction force data. These sensors can capture detailed and dynamic information about body movements and forces involved in walking to analyze gait cycles and parameters. There are 13 studies ($7.7\%$) focused on vision-based devices such as RGB cameras and infrared cameras to extract gait patterns from videos. A same number of studies (13 articles) combine both sensor-based and vision-based devices to collect multi-modal gait data for learning.
Fig. 8e further shows the distribution of sensor types used in 142 sensor-based studies: The force sensors ($e.g.,$ force/pressure sensitive resistors) and IMUs are the two most widely-used sensor types, occupying $55.6\%$ (79 studies) and $31.0\%$ (44 studies) respectively. For other sensors (19 studies), there are 5 studies combining both FS and IMUs as gait data sources, while 5 studies utilize gait analysis system (GAS) ($e.g.,$ mats, walkways, treadmills). The low-cost and non-invasive motion capture device, Kinect [118, 87], is also used as both a sensor and a vision-based device for gait data collection (3 studies). We provide a further discussion on exploiting sensor-based ($e.g.,$ Kinect-based) 3D skeleton data for NDs diagnosis in Sec. 7.5.
@@ 3.2.6 Distribution of AI Models
As shown in Fig. 8a, in the past decade, only 20 articles ($11.8\%$) propose novel A-DL models for NDs diagnosis from gait, while most of the studies use conventional AI methods such as C-ML and C-DL with 105 ($62.1\%$) and 44 articles ($26.0\%$) respectively. However, most conventional AI models possess limited performance on large-scale medical data, thereby underscoring the challenge and urgency of developing more advanced AI models in this field. Such development requires more efforts and cooperation from AI scientists and healthcare professionals in the future.
@@ 3.2.7 Distribution of Sample Sizes
As shown in Fig. 8c, 141 papers ($83.4\%$) conduct their experiments with more than 30 samples ($i.e.,$ number of patients plus number of controls), and the number of papers that possess small-size samples ($10\leq n\leq 30$) only occupies $16.6\%$ (28 papers). This suggests that most AI models require a relatively larger size of samples and gait data to obtain satisfactory performance, otherwise it might be insufficient for obtaining reliable diagnosis. This also aligns with the recent advancements of medical deep learning based AI models that necessitate massive medical labeled data.
@@ 4.1 Quality Evaluation Criterion
To comprehensively measure the quality of each study, we empirically grade the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples at three levels with the score 1, 2, or 3. For novelty (defined as “risk-evaluated novelty”) scores, we not only measure how advance the AI architecture is, but also empirically consider the risk of model over-fitting in real-world application scenarios ($i.e.,$ learning on different sizes of data). For example, C-ML models can be advantageous with higher interpretability and reliability on small datasets, and it may even outperform deep learning methods that could over-fit in some cases [119]. We sum different kinds scores and divide them by the maximum score to get a normalized quality score $S_{Q}$ ranging from 0 to 1. $S_{Q}$ is computed for each study, which are defined as follows:
  • Score ($S_{n}$) for Risk-Evaluated Novelty of AI Models
@@
      We first compute the basic score to assess the novelty level of the AI architecture:
      • Basic Score $\overline{S}_{n}=1$: The main proposed method belongs to or contains only C-ML models.
      • Basic Score $\overline{S}_{n}=2$: The main proposed method belongs to or contains C-DL models.
      • Basic Score $\overline{S}_{n}=3$: The main proposed method belongs to or contains A-DL models.
@@
      We then compute the risk score of over-fitting based on the data size ($i.e.,$ size of subjects):
      • Risk Score $\hat{S}_{n}=1$: The deep learning models (C-DL and A-DL) are learned on the small-size dataset with less than 30 subjects.
      • Risk Score $\hat{S}_{n}=0$:The deep learning models (C-DL and A-DL) are learned on the large-size dataset with more than 100 subjects, or medium-size dataset with more than 30 and less than 100 subjects.
      • Risk Score $\hat{S}_{n}=-1$: The C-ML models are learned on the small-size data with less than 30 subjects, or medium-size dataset with more than 30 and less than 100 subjects.
@@
      Finally, the risk-evaluated novelty score is computed by subtracting the risk score from the basic score, $i.e.$, $S_{n}=\overline{S}_{n}-\hat{S}_{n}$.
@@
Table 1: Number of studies related to different NDs (PD, HD, MS, AD, ALS) and different model types (C-ML, C-DL, A-DL). “Combinatorial” denotes studies that simultaneously contain diagnosis of multiple NDs including PD, ALS, and HD. The numbers of studies under different quality score thresholds $S_{Q}>0.5,0.6,0.7$ are also presented.
       Type        C-ML        C-DL        A-DL        Total        $\boldsymbol{S_{Q}>0.5}$        $\boldsymbol{S_{Q}>0.6}$        $\boldsymbol{S_{Q}>0.7}$
       PD        78        34        14        126        112        56        43
@@
       11        7        3        21        21        12        10
       HD        6        1        0        7        6        2        1
       MS        4        1        1        6        2        2        1
       AD        3        1        2        6        6        4        2
       ALS        3        0        0        3        1        1        0
       Total        105        44        20        169        148        77        57
(a) PD Studies with Different Thresholds
(b) Combinatorial Studies with Different Thresholds
Figure 10: Comparison of numbers of selected PD studies (Fig. (a)) and combinatorial studies (Fig. (b)) when setting different quality score thresholds ($S_{Q}$=0.4, 0.5, 0.6, 0.7).
@@ 4.2 Parkinson’s Disease (PD)
As shown in Table 2, 3, 4, 5 and 6, there are 126 studies that focus on PD diagnosis based on human gait. It is worth mentioning that there are 21 studies, each of which simultaneously includes the diagnosis of PD and other NDs, as shown in Table 6 (referred to as “Combinatorial Diagnosis of Different NDs”). We separately discuss these studies in Sec. 4.3. Based on our survey, there are 78, 34, and 14 PD studies using C-ML, C-DL, and A-DL models , respectively. It should be noted that there are a large number of PD studies (up to 126 studies). Since the studies with high-quality AI models and diagnostic results are the focus of our survey, we first categorize all studies by three AI model types (C-ML, C-DL, A-DL), and then select representative studies with quality score $S_{Q}>0.6$ for content summary and analysis. It is worth noting that we empirically set the quality score threshold to 0.6 because it can select more representative ($e.g.,$ top 40% to 50%) high-quality studies than setting $S_{Q}>0.5$ (88.9% studies), as shown in Fig. 10.
@@ 4.2.1 C-ML Models for PD Diagnosis
Khorasani et al. [120] (Table 2, ID = P-7) utilized Hidden Markov Models (HMM) with Gaussian Mixtures to classify gait data and differentiate PD patients from healthy subjects. Using stride interval as the distinguishing feature, the HMM method achieved an accuracy rate of 90.3%. In their study, the authors, Tupa et al. [85] (ID = P-10). presented a methodology using Microsoft Kinect sensors to detect gait attributes and recognize gait of PD patients. The leg length, stride length, and gait velocity were extracted to achieve high-accuracy classification using neural networks (NN). This study suggested the potential of using Kinect sensors for gait analysis and disorder recognition. Zeng et al. [121] (ID = P-14) proposed a radial basis function neural networks (RBF-NN) based method for classifying PD patients and healthy individuals using gait characteristics derived from vertical ground reaction forces, achieving an over 90% accuracy. Ertuugrul et al. [122] (ID = P-16) focused on a technique called Shifted One-Dimensional Local Binary Patterns (Shifted 1D-LBP) to track local changes in time series data. Combining with machine learning model, this work demonstrated Shifted 1D-LBP’s potential in real-time NDs detection applications. The study of Alam et al. [62] (ID = P-20) utilized the vGRF data to extract gait features and feeded them into the SVM classifier to obtain a good accuracy result of 93.6%. Joshi et al. [123] (ID = P-21) explored the use of wavelet analysis combined with SVM for the identification of PD based on spatio-temporal gait variables, achieving a classification accuracy of 90.32%. The study demonstrated the potential of using wavelet analysis as an efficient method for classifying PD and healthy subjects. Rovini et al. [124] (ID = P-30) utilized a wearable inertial device to collect motor data from healthy subjects, individuals with idiopathic hyposmia (IH), and PD patients during lower limb assessments, and developed a machine learning-based system with SVM, RF, NB to classify the three groups. In their work , Aich et al. [125] (ID = P-32) quantified gait parameters objectively using wearable accelerometers, and compare them with a motion capture system to automatically discriminate patients with PD using machine learning models such as SVM, KNN, and DT. The proposed approach achieves an accuracy of approximately 89.1% and suggests that the wearable accelerometer-based system is suitable for assessing and monitoring PD and Freezing of gait (FoG) in real-life scenarios. Elden et al. [126] (ID = P-38) emphasized the benefit of utilizing a hybrid of linear and non-linear features extracted from human gait time series data to improve the NDs classification system which employed SVM with a radial basis kernel function. Khourly et al. [127] (ID = P-42) leveraged different classic machine learning methods (KNN, DT, RF, NB, SVM, K-Means) to identify PD from vGRFs collected from gait cycles. The structure of PD recognition process from data acquisition to performance evaluation was also systematically presented in this study. Andrei et al. [128] (ID = P-47) devised an automated system with force sensors to extract gait parameters and then used these parameters as input to the SVM model to diagnosis PD. Zeng et al. [129] (ID = P-50) proposed a novel method for classifying gait patterns between patients with PD and healthy controls. This method reconstructed the phase space of vGRF and decomposed the gait dynamics using empirical mode decomposition, which were fed to NN for classification with accuracy of 98.8%. The effectiveness of Extreme Gradient Boosting (XGBoost) and Artificial Neural Network (ANN) models for monitoring the gait of PD patients has been explored, achieving an accuracy of 41.0% (Hughes et al. [130]). The study by Som et al. [131] (ID = P-60) explored the use of single-sensor accelerometer data from healthy subjects to extract features relevant to multi-sensor accelerometer gait data for PD classification. They used a pre-trained convolutional autoencoder as the source model and a simple multi-layer perceptron as the target model to classify PD and achieved a 68.92% accuracy. A machine learning based gait classification system with different models, such as SVM and RT, was presented to assist in diagnosing PD form vGRF (Balaji et al. [132] (ID = P-65)). Balaji et al. also [133] (ID = P-71) extracted spatio-temporal gait features such as stride time and stance time from vGRF and exploits different classic machine learning models (KNN, NB, SVM) to classify health control and PD subjects with an accuracy of 88.5-98.8%. In the work by Ghaderyan et al. [134] (ID = P-81), a time-varying singular value decomposition method was designed to extract the most useful and informative part of the vGRF signal data, and then sparse non-negative least squares (NNLS) approach was leveraged to classify PD and HCs, achieving 97.2% accuracy. Mehra et al. [135] (ID = P-87) utilized RF to identify PD patients from vGRF and achieves up to 98.8% accuracy. The SVM model was used in [136] (ID = P-111) to classify PD and HC gait from vGRF data, with an accuracy of 98.2%. Johnson et al. [137] (ID = P-125, $S_{Q}=0.88$) addressed early-stage PD monitoring and diagnostic challenges using consumer-grade wearable devices and sensor data from the WATCH-PD study. Data were gathered over a year from early-stage PD patients (n = 82) and age-matched controls (n = 50). Leveraging disease- and behavior-relevant feature engineering with multivariate machine learning, the authors developed a PD screening tool based on a random forest model. Achieving 92.3% accuracy, 90.0% sensitivity, and 100% specificity (AUC = 0.92), this approach underscores the potential of remotely monitored wearable devices for population-wide PD screening and monitoring.